A 33.3-g ice cube at 0 °C is added to 126 g of water in a 62.5-g aluminum cup. The cup and the water have an initial temperature of 24.3 °C. Calculate the equilibrium temperature of the cup and its contents. Do not enter unit. Calculate the equilibrium temperature of the cup and its contents if the aluminum cup is replaced with one of equal mass made from copper. Do not enter unit.
Answer #1

This question is based on the concept of heat balance I.e.

Heat gained by one substance is equal to heat lost by another substance in case of mixing.

The detailed solution is described below

Let w goth ihs basics fast mi = 33.3g = 0.0333 K2(@0C) mw = 126 a dc) 0,126 K 2 ti 24.3 e Heat taken ice 0.0625 * goo) (24.3-t) + 0.0333 (334,000 +4186 t ラ(0 . I 26 *4186 ウ111 1 8357-523.(86土ニ 11,122-2+ 33.35 38t

hen alumniun Cur is Yeflaced フ@.126 418( + 0.0625米3R5 )Ca.3-5 L34s1-YL, 551-4325七=11,122.2 +139.393 六= 3.2.58e Ceydi b ,ten@) → hanks